OpenStudy (anonymous):

Convert to higher terms 8/9 = 96/?

OpenStudy (anonymous):

\[12 \times 9 = 108 \]

jhonyy9 (jhonyy9):

because 8*12 =96

OpenStudy (amistre64):

8*n= 96 --- -- 9*n = X

jhonyy9 (jhonyy9):

so 9*12=108

OpenStudy (anonymous):

thank u